Using the number 1,2,3,4,5 and 6 as the element of the population



a. find the mean of the sample of size 3 without replacement


b.construct the sampling distribution of the sample mean



I. The mean


II.The variance


III. The standard deviation of the sampling of the sample means

1
Expert's answer
2022-02-09T13:05:15-0500

The number of possible samples is "\binom{N}{n}=\binom{6}{3}=20".

The possible samples with their means are given below.

"a)"

The sample means are derived from the formula,

"\bar x_i={\sum(x_i)\over3}"

Sample mean

(1,2,3) 2

(1,2,4) 2.33

(1,2,5) 2.67

(1,2,6) 3

(1,3,4) 2.67

(1,3,5) 3

(1,3,6) 3.33

(1,4,5) 3.33

(1,4,6) 3.67

(1,5,6) 4

(2,3,4) 3

(2,3,5) 3.33

(2,3,6) 3.67

(2,4,5) 3.67

(2,4,6) 4

(2,5,6) 4.33

(3,4,5) 4

(3,4,6) 4.33

(3,5,6) 4.67

(4,5,6) 5


"b)"

The sampling distribution is,

"\bar x_i" 2 2.33 2.67 3 3.33 3.67 4 4.33 4.67 5

"p(\bar x _i)" "{1\over20}" "{1\over20}" "{2\over20}" "{3\over20}" "{3\over20}" "{3\over20}" "{3\over20}" "{2\over10}" "{1\over20}" "{1\over20}"



"c)"

The mean of the sampling distribution of the means is the population mean given as,

"\mu_{\bar x_i}=\mu={1+2+3+4+5+6\over6}={21\over6}=3.5"


"d)"

To find the variance of the sampling distribution of the means we first determine the the population variance "\sigma^2", given as,

"\sigma^2={\sum (x^2)-{({\sum (x)})^2\over N }\over N}={91-73.5\over6}=2.91667"

Now, the variance of the sampling distribution of the means is,

"\sigma^2_{\bar x_i}={\sigma^2\over n}={2.91667\over 3}=0.9722"


"e)"

The standard deviation of the sampling of the sample means is,

"\sigma_{\bar x_i}=\sqrt{\sigma^2_{\bar x_i}}=\sqrt{0.9722}=0.9860"